# UnsafeBench: Benchmarking Image Safety Classifiers on Real-World and AI-Generated Images

### v2

@@ Introduction
Counteracting AI-Generated Unsafe Images. Since text-to-image models like Stable Diffusion gained popularity in 2022, concerns have been raised regarding their risks of generating realistic unsafe images. Plenty of studies [59, 52, 70, 71, 27] focus on assessing these models’ risks. Schramowski et al. [59] take the first step in estimating the probability of Stable Diffusion in generating unsafe images when providing harmful prompts. Qu et al. [52] adopt a similar approach and find that text-to-image models are prone to generate sexually explicit, violent, disturbing, hateful, and political images. Other researchers investigate the proactive generation of unsafe images from text-to-image models through various attacks, such as data poisoning attacks [70] and adversarial examples [71, 27]. To mitigate the risks, another line of research focuses on enhancing safety measures. For example, Schramowski et al. [59] propose safe latent diffusion, which steers the generated images away from a list of unsafe concepts during the generation process. Guo et al. [38] takes the first step in using VLMs and chain-of-thought to identify unsafe images from user-generated content in games. All the above works rely on existing image safety classifiers or VLMs to identify the unsafe images generated by text-to-image models. However, since these classifiers are mostly trained on real-world images, it is unclear how effectively they generalize to AI-generated images. Our benchmarking framework, UnsafeBench, investigates their ability to generalize to AI-generated unsafe images and explores the distinct characteristics of those images.
Conclusion & Discussion
@@
We establish UnsafeBench, a benchmarking framework that comprehensively evaluates the effectiveness and robustness of image safety classifiers on both real-world and AI-generated images. We construct a large image safety dataset of 10K manually annotated images and evaluate five conventional classifiers and three VLMs. Our analysis reveals the drawbacks of these classifiers and the unique challenges posed by AI-generated images to both their effectiveness and robustness. We therefore introduce an improved image moderation tool, PerspectiveVision, that accounts for both real-world and AI-generated images. Our work has the following implications for the research community and platform moderators.
Importance of UnsafeBench Dataset. We provide the first large-scale image safety dataset with 11 categories of unsafe content and two sources. Annotated images in this dataset are equipped with three “labels”: 1) safe/unsafe, 2) specific category, and 3) sourced from the real world or AI. These attributes greatly improve the dataset’s utility, enabling future studies to investigate unsafe images in a multifaceted manner. We hope the dataset serves as a solid foundation for future image safety research.
Challenges Posed by AI-Generated Images. In the era of generative AI, the influence of AI-generated content like unsafe imagery and deepfakes can be complex and multifaceted to existing models. Our analysis reveals that they do not only reduce the effectiveness of these models but also render them more vulnerable to adversarial attacks than real-world images. This is because, although current AI-generated images are realistic, there remains a distribution shift from real-world images in image quality statistics (noise level, entropy, etc.), styles, and layouts. For platform moderators and model developers, it is essential to include AI-generated unsafe images in the training data to mitigate these negative impacts.
Utility of PerspectiveVision. PerspectiveVision is built to address the drawbacks of existing image safety classifiers, including imbalanced performance across categories, state-of-the-art model (GPT-4V) being closed-source, and degraded performances on AI-generated images. It is also the first attempt to build an image moderation tool that covers both the wide spectrum of unsafe content and unsafe images from different sources. The research community and platform moderators can directly apply it for various moderation tasks, such as reporting online unsafe images and evaluating the risks of new text-to-image models.
@@
Limitations. Our work has limitations. First, the UnsafeBench dataset is annotated by only three authors, with the final label determined by majority vote. We did not rely on crowdsourcing workers for two reasons: 1) annotation requires expert knowledge in the image safety domain, such as identifying hate symbols, which cannot be guaranteed and would require specific training for crowdsourcing workers; 2) due to ethical considerations, we aim to minimize the number of annotators exposed to such unsafe content and prevent them from exposing it to third parties. Second, images sourced from Lexica are mostly generated by a single text-to-image model, Stable Diffusion. We will expand our dataset with images generated by other text-to-image models and routinely update it. Third, we only evaluate the generalizability of PerspectiveVision on a limited range of unsafe categories due to the lack of labeled datasets. We will continue expanding the evaluation with more datasets and improving the model in future work.
@@ Appendix A Appendix
Q16 [60]. Q16 is a binary image classifier that predicts a given image as morally positive or negative. [^6] https://github.com/ml-research/Q16. It first extracts the image embedding with the CLIP [54] image encoder and calculates the cosine similarity between the image embedding and two soft prompts (optimized text embeddings), one representing the positive class and the other being the negative class. Q16 then assigns the image to the more relevant class. To obtain the optimal soft prompts, the Q16 developer starts with two normal prompts, e.g., “This image is positive/negative,” and iteratively updates their text embeddings to maximize the cosine similarity values on a labeled image dataset, SMID [29]. The Socio-Moral Image Database (SMID) includes 2,941 morally positive and negative images, covering concepts including harm, inequality, degradation, deception, etc. Q16 has been widely used in evaluating the safety/appropriateness of images [60, 62, 59, 44, 73, 35, 28, 66, 27], including identifying unsafe images sourced from LAION-5B [62] and text-to-image models [59, 44, 73, 35, 48], as well as testing the reliability of safety mechanism built in text-to-image models [28, 66, 27].
@@
MultiHeaded Safety Classifier (MultiHeaded) [53]. This classifier is also a CLIP-based model with five linear classification heads, detecting sexually explicit, violent, shocking, hateful, and political images. [^7] https://github.com/YitingQu/unsafe-diffusion. The classifier is trained on 800 labeled images generated by text-to-image models like Stable Diffusion [53]. It has been employed to estimate the cleanliness (percentage of unsafe images) of large public datasets [53] such as LAION-400M [63], LAION-2B [6], LAION-aesthetics [61], and YFCC15M [64].
@@
Stable Diffusion Filter (SD_Filter) [55]. This filter is built in Stable Diffusion to prevent generating explicit images. [^8] https://huggingface.co/CompVis/stable-diffusion-safety-checker. It relies on the CLIP image encoder to assess image embeddings, then measures the distance between these embeddings and text embeddings of 20 sensitive concepts [23], such as “sexual,” “nude,” and “sex.” During the image generation process, the filter compares the embedding distance between the generated image with a set of pre-defined thresholds. If the image is too close to any sensitive concepts, the filter will predict it as unsafe and block it from users.
@@
NSFW_Detector [14]. The NSFW_Detector is also a classification model with the CLIP image encoder as the backbone and a multilayer perceptron (MLP) as the classification head. [^9] https://github.com/LAION-AI/CLIP-based-NSFW-Detector. It mainly detects sexual and nudity images. LAION-AI [8] uses this model to assess the NSFW score for images in their large-scale datasets including LAION-400M [63] and LAION-5B [7].
@@
NudeNet [15]. NudeNet is a lightweight nudity detection tool and consists of a detector that detects the sexually explicit areas in images and a classifier that directly classifies them as explicit images or not. In this study, we use the binary NudeNet classification model. [^10] https://pypi.org/project/nudenet/. It is a classifier with Xception as the backbone, trained on roughly 160K automatically-labeled images [16]. It has been extensively applied in detecting sexually explicit images  [59, 68, 48, 35, 73] and is often combined with Q16 for identifying general unsafe content [59, 35, 48, 73].
@@
LLaVA [45]. LLaVA is an open-source visual language model that can answer questions based on the user’s provided image and prompt. It comprehends the image with the CLIP image encoder and understands the user’s prompts with a large language model, Vicuna [24]. LLaVA is trained on the LAION-CC-SBU dataset and instruction dataset generated by GPT4V [45]. We use the llava-v1.5-7b checkpoint. [^11] https://huggingface.co/liuhaotian/llava-v1.5-7b.
@@
Adversarial examples are an important tool to evaluate a classifier’s robustness. We introduce three widely used gradient-based methods to create adversarial examples, FGSM [37], PGD [47], and DeepFool [50].
@@
PGD (Projected Gradient Descent) [47]. PGD is an iterative version of FGSM and updates adversarial examples step by step. To obtain the adversarial example at time $t+1$, it first applies perturbations on the adversarial examples at time $t$ with the step size of $\alpha$. It then projects them back to the $\epsilon$-ball space (to constrain the perturbations).
@@
Experimental Setup
  • LoRA Fine-Tuning. We build the instruction dataset using the prompt template shown in Figure 10. Specifically, we construct a pair of instructional data points for each unsafe image, which we call the positive example and the negative example. In the positive example, given the unsafe image, we first identify the unsafe category it violates and then randomly append K (from 1 to 10) other unsafe categories to create a customized image safety taxonomy. The safety label is still recorded as unsafe. Meanwhile, in the negative example, we randomly select K unsafe categories that the given unsafe image does not violate, and then we record the safety label as safe. This approach ensures that VLMs learn the safety of images rather than simply memorizing the output format. We then fine-tune LLaVA on this dataset with its original training loss (CrossEntropy loss) and optimizer (AdamW), following the setup in [45]. The batch size is 16, and the learning rate is 2e-4. We set the training epoch to 1.

### v3

@@ A Case Study on Artistic Representation and Grid Layout
Counteracting AI-Generated Unsafe Images. Since text-to-image models like Stable Diffusion gained popularity in 2022, concerns have been raised regarding their risks of generating realistic unsafe images. Plenty of studies [59, 52, 67, 68, 26, 39] focus on assessing these models’ risks. Schramowski et al. [59] take the first step in estimating the probability of Stable Diffusion in generating unsafe images when providing harmful prompts. Qu et al. [52] adopt a similar approach and find that text-to-image models are prone to generate sexually explicit, violent, disturbing, hateful, and political images. Other researchers investigate the proactive generation of unsafe images from text-to-image models through various attacks, such as data poisoning attacks [67] and adversarial examples [68, 26]. To mitigate the risks, another line of research focuses on enhancing safety measures. For example, Schramowski et al. [59] propose safe latent diffusion, which steers the generated images away from a list of unsafe concepts during the generation process. Guo et al. [37] takes the first step in using VLMs and chain-of-thought to identify unsafe images from user-generated content in games. All the above works rely on existing image safety classifiers or VLMs to identify the unsafe images generated by text-to-image models. However, since these classifiers are mostly trained on real-world images, it is unclear how effectively they generalize to AI-generated images. Our benchmarking framework, UnsafeBench, investigates their ability to generalize to AI-generated unsafe images and explores AI specific characteristics.
Conclusion
@@
We establish UnsafeBench, a benchmarking framework that comprehensively evaluates the effectiveness and robustness of image safety classifiers on both real-world and AI-generated images. We construct a large image safety dataset of 10K manually annotated images and evaluate five conventional classifiers and three VLMs. Our evaluation reveals how AI-generated unsafe images pose a challenge to existing classifiers in terms of both effectiveness and robustness. We further introduce the image moderation tool, PerspectiveVision, to capture the distribution shift from AI images in image quality statistics (noise level, etc.), styles, and layouts.
@@
Limitations. Our work has limitations. First, the UnsafeBench images (both safe and unsafe images) are collected using the same set of unsafe keywords, which makes them more challenging to classify for image safety classifiers than using irrelevant safe keywords (e.g., cats, dogs). However, we intentionally designed this curation process to collect challenging borderline examples, such that the classifiers are expected to distinguish between truly unsafe content and similar but acceptable content. Second, the UnsafeBench dataset is annotated by three experts in the research team. We did not rely on crowdsourcing workers for two reasons: 1) annotation requires expert knowledge in the image safety domain; and 2) due to ethical considerations, we aimed to prevent unsafe content from being exposed to third parties. Third, images sourced from Lexica are mostly generated by a single text-to-image model, Stable Diffusion. We will expand our dataset with images generated by other text-to-image models and routinely update it.
Acknowledgements
We thank all anonymous reviewers for their constructive suggestions. This work is partially funded by the European Health and Digital Executive Agency (HADEA) within the project “Understanding the individual host response against Hepatitis D Virus to develop a personalized approach for the management of hepatitis D” (DSolve, grant agreement number 101057917) and the BMBF with the project “Repräsentative, synthetische Gesundheitsdaten mit starken Privatsphärengarantien” (PriSyn, 16KISAO29K).
@@ Appendix A Appendix
Classifier Backbone Training Paradigm Training Dataset
Q16 CLIP Prompt Learning SMID [28] (R)
MultiHeaded CLIP Linear Probing MultiHeaded-D [53] (A)
SD_Filter CLIP - -
NSFW_Detector CLIP Linear Probing LAION-400M Subset (R)
NudeNet Xception Fully Supervised NudeNet-D [15] (R)
Q16 [60]. Q16 is a binary image classifier that predicts a given image as morally positive or negative. [^6] https://github.com/ml-research/Q16. It is trained on a labeled image dataset, SMID [28], which includes 2,941 morally positive and negative images, covering concepts including harm, inequality, degradation, deception, etc.
@@
MultiHeaded Safety Classifier (MultiHeaded) [53]. This classifier is also a CLIP-based model with five linear classification heads, detecting sexually explicit, violent, shocking, hateful, and political images. [^7] https://github.com/YitingQu/unsafe-diffusion. The classifier is trained on 800 labeled images generated by text-to-image models like Stable Diffusion [53].
@@
Stable Diffusion Filter (SD_Filter) [55]. This filter is built in Stable Diffusion to prevent generating explicit images. [^8] https://huggingface.co/CompVis/stable-diffusion-safety-checker. It relies on the CLIP image encoder to assess image embeddings, then measures the distance between these embeddings and text embeddings of 20 sensitive concepts [22], such as “sexual,” “nude,” and “sex.
@@
NSFW_Detector [13]. The NSFW_Detector is also a classification model with the CLIP image encoder as the backbone and a multilayer perceptron (MLP) as the classification head. [^9] https://github.com/LAION-AI/CLIP-based-NSFW-Detector. It mainly detects sexual and nudity images.
@@
NudeNet [14]. NudeNet is a lightweight nudity detection tool and consists of a detector that detects the sexually explicit areas in images and a classifier that directly classifies them as explicit images or not. We use the binary NudeNet classification model. [^10] https://pypi.org/project/nudenet/.
Details of the RoBERTa Classifier
Table 8: Reliability of the RoBERTa classifier.
VLM Accuracy Precision Recall F1-Score
LLaVA 1.000 1.000 1.000 1.000
InstructBLIP 0.980 0.981 0.980 0.978
GPT-4V 0.995 0.995 0.995 0.995
Overall 0.992 0.992 0.992 0.991
To convert VLM-generated responses into standard classes, we fine-tune a language model, RoBERTa [46], to categorize LLM outputs into one of three classes: safe, unsafe, or uncertain. Since most responses already contain indicative class terms, only a few consist of lengthy sentences. To this end, we annotate a small set of 10–20 unique responses per class for each VLM and train the classifier for 10 epochs. For example, for the safe class, we use outputs such as “safe,” “the image is safe,” “the image is not considered unsafe as…” as representative examples and fine-tune RoBERTa in a supervised manner. To validate the reliability of this classifier, we randomly sample 600 VLM-generated responses (200 from each VLM) and manually classify them into safe, unsafe, and uncertain classes. We then compare this ground truth with the predictions from the RoBERTa classifier and compute the metrics shown in Table 8. The evaluation confirms the high reliability of the classifier.
@@
LLaVA [45]. LLaVA is an open-source visual language model that can answer questions based on the user’s provided image and prompt. It comprehends the image with the CLIP image encoder and understands the user’s prompts with a large language model, Vicuna [23]. We use the llava-v1.5-7b checkpoint. [^11] https://huggingface.co/liuhaotian/llava-v1.5-7b.
@@
PGD (Projected Gradient Descent) [47]. PGD is an iterative version of FGSM and updates adversarial examples step by step. To obtain the adversarial example at time $t+1$, it first applies perturbations on the adversarial examples at time $t$ with the step size of $\alpha$. It then projects them back to the $\epsilon$-ball space.
@@
Experimental Setup
(a) The Sexual Category
(b) The Violence Category
Figure 10: Optimal K for the Sexual and Violence category.
  • LoRA Fine-Tuning. We build the instruction dataset using the prompt template shown in Figure 9. Specifically, we construct a pair of instructional data points for each unsafe image, which we call the positive example and the negative example. We then fine-tune LLaVA on this dataset with its original training loss (CrossEntropy loss) and optimizer (AdamW), following the setup in [45]. The batch size is 16, and the learning rate is 2e-4. We set the training epoch to 1.
@@
Table 12: F1-Score of PerspectiveVision models and baselines on the UnsafeBench test set.
Type Model Hate Harassment Violence Self-harm Sexual Shocking Illegal Activity Deception Political Health Spam Overall
PerspectiveVision Linear Probing (CLIP) 0.776 0.736 0.767 0.586 0.924 0.894 0.877 0.841 0.907 0.903 0.899 0.859
Prompt Learning (CLIP) 0.437 0.557 0.679 0.337 0.890 0.758 0.765 0.596 0.829 0.635 0.654 0.687
LoRA Fine-tuning (LLaVA) 0.686 0.755 0.760 0.743 0.913 0.861 0.889 0.780 0.878 0.909 0.855 0.844
Classifier Ensemble Q16_NudeNet 0.543 0.454 0.619 0.368 0.704 0.735 0.591 0.679 0.515 0.538 0.219 0.585
Q16_NSFW_Detector 0.571 0.503 0.646 0.400 0.761 0.765 0.576 0.693 0.504 0.500 0.147 0.606
Q16_SD_Filter 0.473 0.430 0.610 0.396 0.849 0.710 0.566 0.648 0.507 0.449 0.205 0.595
Q16_MultiHeaded 0.550 0.495 0.626 0.421 0.788 0.773 0.580 0.660 0.713 0.535 0.216 0.635
Q16_MultiHeaded_NudeNet 0.524 0.446 0.608 0.359 0.826 0.738 0.588 0.649 0.730 0.566 0.247 0.625
All Conventional Classifiers 0.444 0.469 0.589 0.331 0.861 0.700 0.571 0.621 0.723 0.512 0.260 0.607
Commercial Google’s SafeSearch 0.095 0.423 0.519 0.308 0.826 0.623 0.307 0.200 0.589 0.442 0.279 0.515
Microsoft’s Filter 0.341 0.310 0.642 0.304 0.854 0.435 0.298 0.172 0.125 0.299 0.127 0.475
GPT-4V 0.439 0.626 0.747 0.500 0.886 0.819 0.848 0.636 0.780 0.500 0.549 0.717